1. Find the theta notation of the following:

a. 6n ³+ 12n ² + 1, for nà ƒ ƒ ¢ ‰ ¥1

b. 3n ²+ 2n log2 n, for nà ƒ ƒ ¢ ‰ ¥1

2. Find the theta notation for the number of times the statement x:=x + 1 is executed.

a. for i:=1 to 2n do x=x+1

b. for i;=1 to 2n do for i;=1 to n do X:=x +1
1
Expert's answer
2021-10-20T18:02:17-0400

1.

Big-Theta(Θ) notation gives bound for a function f(n) to within a constant factor.

We write f(n) = Θ(g(n)), If there are positive constants n0 and c1 and csuch that, to the right of n0 the f(n) always lies between c1g(n) and c2g(n) inclusive.

Θ(g(n)) = {f(n) : There exist positive constant c1, c2 and n0 such that 0 ≤ c1 g(n) ≤ f(n) ≤ cg(n), for all n ≥ n0


a.

"c_1n^3\le" 6n ³+ 12n ² + 1"\le c_2n^3"

"f(n)=\theta(n^3)"


b.

"c_1n^2\le" 3n ²+ 2n log2 n"\le c_2n^2"

"f(n)=\theta(n^2)"


2.

a.

"f(n)=2n"

"c_1n\le 2n\le c_2n"

"f(n)=\theta(n)"


b.

"f(n)=2n^2"

"c_1n^2\le 2n^2\le c_2n^2"

"f(n)=\theta(n^2)"